If A = 2i+3j+8k is perpendicular to  B = -4i+4j+yk, then what is the value of ?

Grade:12

3 Answers

Arun
6 years ago
Dear Akshay
 
If A and B are perpendicular then
A. B = 0
(2i + 3j +8k) . (-4i + 4j + yk) = 0
-8 +12 + 8y = 0
8y = -4
y = -1/2
Regards
Arun (askIITians forum expert)
Khimraj
6 years ago
If two vector are perpendicular then their dot product will be zero.
A.B = 0
-8 + 12 +8y = 0.
So y = - 1/2 
Meghna
6 years ago
If two vector are perpendicular so there dot product will be zero.A.B = -8i + 12j +8y k 0 = -8 +12 +8y-8y = +4 Y= -1/2